OpenStudy (anonymous):

express radical 18 + 5 radical 2 in simplest form. how?

OpenStudy (kropot72):

\[\sqrt{18}+5\sqrt{2}\] The way to solve this is to first find the factors of 18. Can you factorise 18?

OpenStudy (anonymous):

Yes doesn't radical 18 end up being 3 radical 2?

OpenStudy (whpalmer4):

it does! and you can add them together to get...

OpenStudy (kropot72):

\[3\sqrt{2}+5\sqrt{2}=?\]

OpenStudy (anonymous):

8 radical 2

OpenStudy (kropot72):

You are correct :)

OpenStudy (anonymous):

Thank you!

OpenStudy (kropot72):

You're welcome :)